What average net force is required to bring a 1500kg car to rest from a speed of 100km/h within a distance of 55m?


1
Expert's answer
2021-05-21T10:44:41-0400

First, let's find the deceleration of the car:


"v^2=v_0^2+2as,""a=\dfrac{v^2-v_0^2}{2s},""a=\dfrac{0-(27.77\ \dfrac{m}{s})^2}{2\cdot55\ m}=-7\ \dfrac{m}{s^2}."


The sign minus means that the car decelerates.

Finally, we can find average net force required to bring the car to rest:


"F_{net}=ma=1500\ kg\cdot(-7\ \dfrac{m}{s^2})=-10500\ N."

The magnitude of the average net force is 10500 N. The sign minus means that the average net force directed in the opposite direction to the motion of the car.